OpenStudy (anonymous):

ln(x+7) + ln(x+3)= ln77

OpenStudy (ghazi):

Ln(A*B)= Ln A+Ln B...therefore Ln(x+7)+Ln(x+3)= Ln[(x+7)(x+3)]=Ln 77 .... (x+7)(x+3)=77...solve for X

OpenStudy (ghazi):

can you do it now?

OpenStudy (anonymous):

yes that was easy thanks!

OpenStudy (ghazi):

:)

mathslover (mathslover):

@yeller33 \[\huge{\color{red}{\cal{WELcome}}\color{blue}{\space \mathbb{TO}}\color{green}{\space \mathsf{Openstudy}}}\]

OpenStudy (anonymous):

thankyou mathslover! its a huge help!